Shift report 20011101 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 01 2001 - 08:13:06 EST

  • Next message: Apache: "Shift report 20011101 08:00-16:00"

    ------------------------------------------------------------
    Date: Thu Nov  1 08:13:06 2001, Shift: 00:00-08:00
    Supervisor: DO, with: 
    Log book#10 pages 80 - 81
    
    Shift Summary:
    ------------------------------------------------------------
    1.Machine status
      At 00:28, the beam died suddenly.I had to warn MCR 
      about it (the guy I called didn't notice). He
      told me they had a cryo problem. At 1:30, new store that 
      lasted around 5.5 hours.
    
    2.Achievements:
      Data runs  trig6  Spectrometer setting
        5607      150k    MRS 45 - 4A   FS 12 - 1/4A 
        5609      110k    MRS 60 - 3A   FS 12 - 1/4A
        5610      103k    idem
        5611      138k    idem 
    
      Pedestal run 5614, TPC pedestal 5612
      DVM run 5613
    
    3.No detector problems at all! :) 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 01 2001 - 08:13:56 EST